OpenStudy (anonymous):

m^2-49/2m=1/2 solve by completing the square

OpenStudy (anonymous):

m^2-49/2m+12.25^2=1/2. i bleieve you can solve from there

OpenStudy (anonymous):

m^2-49/2 m+(49/4)^2=1/2+(49/4)^2